OpenStudy (anonymous):

If 4 is subtracted from twice a number, the result is 12 less than the number. A. 2x – 4 = x + 12 x = 16 B. 2x + 4 = x – 12 x = –16 C. 2x + 4 = x + 12 x = 8 D. 2x – 4 = x – 12 x = –8

OpenStudy (anonymous):

2x-4=x-12 Solve for x

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

2x-4=x-12 2x-x=-12+4 x=-8